# The Adoption and Efficacy of Large Language Models: Evidence From Consumer Complaints in the Financial Industry

### v3

@@ 3 Controlled Lab Experiments: Presentation Effect of LLM
Consistent with our observational data, the complaints modified by ChatGPT (Coherent, Clear, or Professional) were more likely to receive hypothetical monetary compensation (M = 3.52, SD = 1.91) than the unedited (Control) complaints (M = 2.78, SD = 1.78), Welch’s t(517) = 5.78, p < .001, as illustrated in Figure 3 (Panel A) and in line with the results from the same comparison in the preregistered linear fixed effects model (b = 0.74, SE = 0.093, t = 7.96, p < .001). These results corroborate the findings from Section 2.3, confirming that LLM-enhanced complaints significantly increase the likelihood of obtaining relief.
Moreover, the results also revealed that complaints edited to improve on specific linguistic feature (i.e., Clear, Coherent, and Professional complaints) were perceived as generally improved in all other linguistic features as well (see Figure 3, Panel B). For example, complaints edited to be clearer exceeded Control (unedited) complaints not only in clarity (b = 1.53, SE = 0.10, t = 14.94, p < .001), but also in coherence (b = 1.61, SE = 0.10, t = 15.60, p < .001) and professionalism (b = 2.23, SE = 0.11, t = 20.48, p < .001). These results further support the notion that LLM editing improves the overall presentation quality of complaints across multiple linguistic features simultaneously (Niven and Kao, 2019; Manning et al., 2020; Liu et al., 2019). This improvement in presentation quality underscores that enhancements in presentation alone—without substantive content changes—significantly influence the effectiveness of complaints. This finding reinforces the idea that the effectiveness of LLMs observed in our CFPB data might predominantly stem from improvements in presentation.
Figure 3: Experimental Results on the Comparison Between Human Complaints and AI Complaints
Note. Panel A shows the mean compensation rating (and its 95% CI). Panel B shows the simultaneous improvement in all three linguistic features (relative to the unedited, Control complaints) for all three types of edited complaints.
Robustness and Relevance Tests
To verify the robustness of our findings, we conducted two additional experiments. The first of these, with a larger sample size (N = 301), focused solely on compensation likelihood by removing the evaluation of linguistic features, which minimized potential bias. This test reaffirmed our initial results, with edited complaints consistently showing a higher probability of receiving compensation. Detailed results can be found in the Online Appendix, 2D.
The second additional experiment (N = 400) was conducted with participants who had prior work experience in the finance industry, thereby enhancing the external validity of our findings. Participants were asked to evaluate complaints based on a specific presentation feature, either clarity or professionalism (see Online Appendix, 2E). The results replicated those of the previous experiments, confirming that the observed effects are robust in more ecologically valid samples and are consistent even when participants focus on different presentation features. The results from all three experiments suggest that LLMs can enhance the effectiveness of complaints primarily by improving overall presentation quality, thereby catering to the diverse preferences of complaint handlers across different linguistic dimensions.
4 Conclusion and Implications
Large language models are among the most transformative advancements in AI technology. Despite their widespread influence, the degree to which consumers adopt LLMs for interacting with firms, as well as the actual efficacy of LLMs in consumer-firm communication, both remain significant empirical questions. This paper contributes as one of the first empirical studies to explore and document the rapid adoption of LLMs by consumers for drafting complaints to financial firms following the release of ChatGPT. Our analysis of over 1 million complaints from the CFPB database shows a significant and sharp increase in LLM usage. To address concerns regarding endogeneity and non-random selection into LLM usage, we employed instrumental variables, Internet Access and English Proficiency, that influence only the usage of LLMs, without directly affecting the probability of obtaining relief. This analysis using IVs helps us establish a causal relationship between Likely-AI complaints and an increased likelihood of obtaining financial relief. Our research also complements the extensive analysis of CFPB data with experimental validation. Through experiments, we demonstrate that the enhancements in presentation quality facilitated by LLMs are crucial in increasing the effectiveness of consumer communications and the likelihood of achieving financial relief.
The findings from this study have substantial implications for consumer advocacy, firm practices, and regulatory policies. By facilitating access to LLM technologies such as integrating LLM-aided tools on their websites where customers submit complaints, firms can enhance consumer satisfaction, particularly aiding those who may struggle with articulating their complaints effectively. Leveraging LLMs to improve communication allows firms to focus on the actual content of disputes rather than being influenced by presentation style. This approach could uncover previously underexplored consumer segments whose concerns were less vocal or inadequately expressed, leading to more effective customer service and complaint resolution strategies. Ultimately, this fosters stronger relationships between firms and consumers, ensuring that genuine issues are addressed more efficiently.
From a policy perspective, equitable access to LLM technology is crucial to prevent exacerbating existing disparities in consumer outcomes. As our research suggests, LLM usage can significantly enhance the likelihood of obtaining relief from firms, primarily due to the improved presentation style, and clarity provided by these tools. However, without equal access, less tech-savvy or financially constrained consumers may find themselves at a disadvantage. Policymakers should consider strategies to broaden access to LLM tools, such as making them available through public libraries, consumer advocacy groups, or other community resources.
Moreover, regulatory bodies like the CFPB should play a proactive role in ensuring that LLM technologies are accessible to all consumers, particularly those from vulnerable groups. By promoting the integration of these tools across socio-economic divides, regulators can help level the playing field and ensure that all consumers have equal opportunities to effectively voice their concerns, enhancing the overall fairness and effectiveness of financial dispute resolutions.

### v4

@@ Results
Complaints edited with ChatGPT (Coherent, Clear, and Professional complaints) were more likely to receive hypothetical monetary compensation (M = 3.92, SD = 2.04) than the unedited (Control) complaints (M = 3.65, SD = 2.06), t(1503) = 2.27, p = .023, as illustrated in Figure S6. More importantly, the greater likelihood of compensation was also found in the preregistered linear fixed effects models controlling for the effects of complaint contents and participants; see Table 4 in the main text. Specifically, the LLM-edited complaints were more likely to be offered hypothetical monetary compensation than the unedited complaints (b = 0.31, SE = 0.087, t = 3.52, p < .001). Thus, results from Experiment 1 directly support the finding from analyses of the CFPB data that editing complaints with an LLM significantly increases the likelihood of obtaining relief.
Figure S6: Compensation Likelihood for Unedited and LLM-Edited Complaints (Experiment 1 Results)
Note. The figure shows the results from Experiment 1. The left portion shows the mean and SE of the compensation ratings for the 20 unedited complaints (Control) and 60 edited complaints (Clear, Coherent, and Professional combined), while the right portion shows the same results separately for each type of LLM edits.
3B Experiment 2
We conducted Experiment 2 with two goals in mind. First, we sought to replicate the findings from Experiment 1 with participants who have worked in the finance industry. Second, we tested whether the previous results would be replicated also when participants focused on specific presentation features of complaints.
Using Prolific, we recruited 400 individuals who have worked in the finance sector in the past ($\textit{M}_{age}$ = 45; 50.5% men, 49% women, and 0.5% other). We began the study by giving the participants some ideas about the study’s purpose (“We’re interested in learning about how consumer complaints might be handled by financial institutions”) and informing them of why they were recruited (“We sought to recruit participants like you who… have past work experience in the finance sector”). Participants then learned that they would encounter four different consumer complaints and were asked to imagine that they were in charge of handling each of the complaints. The key manipulation in Experiment 2 was asking participants to focus on one linguistic feature—clarity or professionalism—and asking them to use it as the main judgment criterion to decide their likelihood of providing compensation. Specifically, for each complaint, we asked participants to rate it on the focal linguistic feature they were assigned (either clarity or professionalism) using the following 7-point scale: “How clear [professional] is the complaint above? (1 = Not clear [professional] at all, 7 = Extremely clear [professional] ).” Participants then rated how likely they would be to provide compensation for the complaint: “If you were handling the complaint above, how likely would you be to offer monetary compensation to address the consumer’s complaint? (1 = Not likely at all, 7 = Extremely likely).” Participants provided these two ratings for each of the four randomly assigned complaints. They then provided demographic information to complete the study.
Experiment 2 replicated the findings from Experiment 1. Specifically, the complaints edited with ChatGPT (Clear, Coherent, and Professional complaints combined) were more likely to receive hypothetical monetary compensation than the unedited complaints (Control complaints only) in a linear fixed effects model which controlled for fixed effects of participants and complaint contents, b = 1.01, SE = 0.084, t = 12.10, p < .001; see Model 2 in Table S7. When we estimate the same model separately within each of the two conditions (Focus on Clarity or Focus on Professionalism), we obtain similar results; see Models 5 and 6 in Table S7. As in Experiment 1, we also tested alternative models (see Table S7).
Table S7: Results From Experiment 2
Dependent variable: Compensation likelihood
Model
1 2 3 4 5 6
Constant 3.337∗∗∗
Unedited vs. Edited (Dummy variable) 0.970∗∗∗ 1.010∗∗∗ 0.990∗∗∗ 1.011∗∗∗ 1.002∗∗∗
(0.114) (0.084) (0.084) (0.122) (0.115)
Unedited vs. More clear (Dummy variable) 0.987∗∗∗
Unedited vs. More coherent (Dummy variable) 1.090∗∗∗
Unedited vs. More professional (Dummy variable) 0.894∗∗∗
Condition All All All All Focus on Clarity Focus on Professionalism
Complaint Content (20 Different Contents) Fixed Effect N Y Y Y Y Y
Participant Fixed Effect N Y Y Y Y Y
Complaint Presentation Position (1-5) Fixed Effect N N Y Y N N
Observations 1,600 1,600 1,600 1,600 800 800
Note. p $<$ .05; ∗∗p $<$ .01; ∗∗∗p $<$ .001
Experiment 2 replicated the findings from Experiment 1. Specifically, the complaints edited with ChatGPT (Coherent, Clear, and Professional complaints combined) were more likely to receive hypothetical monetary compensation than the unedited complaints (Control complaints only) in a preregistered linear fixed effects model, regardless of whether participants focused on clarity or professionalism of complaints, bs > 1.00, SEs < 0.12, ts > 8.25, ps < .001.
3C Experiment S1
Method
Experiment S1 was a nearly direct replication of Experiment 1. We recruited 210 participants ($\textit{M}_{age}$ = 44; 49.5% men, 50% women, and 0.5% other) from Prolific. The procedure of Experiment S1 matched that of Experiment 1, with a minor difference: For each of the five complaints, after rating their likelihood of offering hypothetical monetary compensation (i.e., the main dependent measure), participants also rated the complaint on clarity, coherence, and professionalism (in that fixed order). Specifically, they answered the following three questions right after the main dependent measure, “How clear [coherent / professional] was the complaint above?” (1 = Not clear [coherent / professional] at all, 7 = Extremely clear [coherent / professional]). Aside from this difference, all other aspects of the procedure were the same as in Experiment 1.
Results
We analyzed data exactly as specified in the preregistration (https://aspredicted.org/MM3_17T). Specifically, we fit a linear fixed effects model with the likelihood of offering (hypothetical) monetary compensation as the dependent variable and whether or not a given complaint was edited (to be more clear, coherent, or professional) or not as the independent variable, controlling for (1) the fixed effects of each of the 20 complaint contents and (2) the fixed effects of each of the 210 individual participants (see Model 2 in Table S8).
Table S8: Results From Experiment S1
Dependent variable: Compensation likelihood
Model
1 2 3 4
Constant 2.779∗∗∗
Unedited vs. Edited (Dummy variable) 0.737∗∗∗ 0.743∗∗∗ 0.742∗∗∗
(0.132) (0.093) (0.093)
Unedited vs. More clear (Dummy variable) 0.810∗∗∗
Unedited vs. More coherent (Dummy variable) 0.720∗∗∗
Unedited vs. More professional (Dummy variable) 0.698∗∗∗
Complaint Content (20 Different Contents) Fixed Effects N Y Y Y
Participant Fixed Effects N Y Y Y
Complaint Presentation Position (1-5) Fixed Effects N N Y Y
Observations 1,050 1,050 1,050 1,050
Note. ∗∗∗p $<$ .001
Consistent with our hypothesis (and as mentioned in the main text), the complaints edited with ChatGPT were more likely to receive (hypothetical) monetary compensation than the unedited complaints, b = 0.743, SE = 0.093, t = 7.96, p < .001 (Model 2 in Table S8). Although not preregistered, we fit another linear fixed effects model, which also controlled for the presentation order of complaints (Model 3 in Table S8). We find the same results, b = 0.742, SE = 0.093, t = 7.94, p < .001. Lastly, we examined which type of edits (“more clear” vs. “more coherent” vs. “more professional”) results in the biggest increase in the likelihood of receiving (hypothetical) monetary compensation by fitting yet another linear fixed effects model (Model 4 in Table S8), adding the three types of edit as the three independent dummy variables (unedited = 0, each type of edit = 1) and removing the previous binary variable of whether a given complaint was edited or not. We find that the complaints edited to be more clear showed the directionally greatest increase in likelihood of monetary compensation (b = 0.810, SE = 0.116), as compared with those edited to be more coherent (b = 0.720, SE = 0.115), or those edited to be more professional (b = 0.698, SE = 0.116); see Model 4 in Table S8.
Figure S7: The figure summarizes the procedure and results of Experiment S1. The 20 original complaints were edited with ChatGPT to be more clear, more coherent, or more professional, which resulted in 20 unedited and 60 edited complaints (20 Control, 20 Clear, 20 Coherent, and 20 Professional complaints). Participants in Experiment 1 were presented with a random set of 5 complaints from this pool of 80, viewing one complaint at a time. For each of the 5 complaints, they imagined handling the complaint and first rated the likelihood of offering monetary compensation in response to the complaint, and then evaluated the complaint’s clarity, coherence, and professionalism.
Appendix 4 Note on the Design of Experiments
For the main experiments, we could have chosen to manipulate linguistic features using a study design in which each participant would be randomly assigned to receive one complaint out of the 80 (Alternative Design), rather than a series of multiple complaints (e.g., 5 complaints in a within-subjects design). However, we had a few justifications for opting for our preregistered, within-subjects study designs instead.
First, presenting multiple complaints as in our design likely led participants to form their own standard for good complaints (which in turn would make them better judges), whereas such a process would be unlikely to occur if participants saw only one complaint out of the 80 (Alternative Design), as is typically done in between-subjects experiments where often a single, rather than multiple, stimulus is presented. By presenting multiple complaints, participants were naturally led to compare complaints to one another, a process through which they likely came to establish their own standards of what makes a good complaint. Such standards then would have made them better judges of complaints, compared to evaluating a single complaint (Alternative Design).
Second, our design more closely matched the real-world setting where evaluators of consumer complaints encounter a wide variety of complaint types and multiple of them, rather than exclusively encountering one specific type of complaint or one specific complaint. Thus, exposing participants to multiple complaints of different types made our study more ecologically valid compared to exposing them to a single complaint (Alternative Design). In sum, our study design likely enhanced the robustness and real-world applicability of our study’s findings.